# Balancing Specialization and Adaptation in a Transforming Scientific Landscape

### 03 Jan 2024 v2

@@ A.3.1 Keyword classification
The average effective amount of topics in documents according to the topic model – measured as $\exp H(\bm{\theta}_{d})$ where $H$ is the Shannon entropy – is 7.5, which is unrealistically high. By performing this filtering and classification procedure, the average effective amount of topics goes down to 3.2 per document, which is much more informative and realistic. This extra step is especially important for short texts such as abstracts, which poorly constrain the latent topic distribution $\bm{\theta}_{d}$.
@@ A.4 Measures of capital
Figure 13 shows the Pearson correlation between these diversity and power measures, as evaluated in the cohort of high-energy physicists. For comparison purposes, an alternative measure of diversity based on Stirling’s index [57], with prior applications to studies of interdisciplinarity [58, 59] is evaluated [^18] The Stirling-based diversity measured is evaluated as: $D_{\text{Stirling}}=1-\sum_{k,k^{\prime}}d_{kk^{\prime}}I_{ak}I_{ak^{\prime}}$ (15) Where $d_{kk^{\prime}}$ is the fraction of scientists who have more expertise than average in both $k$ and $k^{\prime}$ among those that have expertise in one or the other (i.e., a similarity matrix); see [58, 59]. A measure of brokerage is also considered [^19] We evaluated brokerage as the amount of pairs of scientists that have collaborated with a given physicist while having no common collaborator except for this physicist. This effectively measures the extent to which this physicist connects otherwise disconnected scientists..
@@ A.6 Learning costs and optimal transport
The MCMC algorithm from [53] is run on 500,000 iterations. The mean posterior values of $C_{k,k^{\prime}}$ are shown in Figure 13(a), as a function of the linguistic gap from $k$ to $k^{\prime}$, $\Delta_{kk^{\prime}}=H(\bm{\varphi_{k^{\prime}}}+\bm{\varphi_{k}})-H(\bm{\varphi_{k}})$ (where $H$ is the Shannon entropy and $\bm{\varphi_{k}},\bm{\varphi_{k^{\prime}}}$ are the vocabulary distributions of $k$ and $k^{\prime}$, such that $\varphi_{kw}$ is the frequency of the keyword $w$ in the topic $k$). For replication purposes, the same procedure is performed on a different dataset – the ACL anthology corpus of computational linguistics research [15]. The correlation is smaller for the ACL cohort ($R=0.82$ instead of $0.96$). The difference could be explained by a number of reasons: a) the smaller sample size of the ACL cohort ($N=578$ instead of $N=2195$); b) differences in the quality of the topic models c) the cognitive scale (on which the overlap in vocabulary between research areas depends) is not commensurable across the two corpora, and finally d) the relationship does not hold as well for fields undergoing major disruptions.
(a) High-energy physics
(b) Computational Linguistics
Figure 14: Cost of shifting a unit of cohort’s research attention from $k$ to $k^{\prime}$ as a function of the fraction of physicists with expertise in $k^{\prime}$ among those with expertise in $k$ ($\nu_{kk^{\prime}}$).
@@ A.9 Replication corpus
For purposes of replication, certain analyses have been reproduced on the ACL anthology corpus of Computational Linguistics research [15]. A comparison with the HEP corpus in shown in Table 7. The ACL corpus is smaller, presumably due to a lack of exhaustivity. The requirement on the quantity of publications has been lowered (three per time period instead of five) in order to maintain a reasonable size of cohort. The choice of time periods is arbitrary (i.e. it was not informed by the context). Only 14 topics were deemed to carry scientific content valid upon manual inspection.

### 14 Jan 2024 v3

@@ A.3.1 Keyword classification
The average effective amount of topics in documents according to the topic model – measured as $\exp H(\bm{\theta}_{d})$ where $H$ is the Shannon entropy – is 7.5, which is unrealistically high. The filtering and classification procedure reduces the average effective amount of topics to 3.2 per document, which is much more informative. This procedure is especially important for short texts such as abstracts, which poorly constrain the latent topic distribution $\bm{\theta}_{d}$.
@@ A.4 Measures of capital
Figure 13 shows the Pearson correlation between these diversity and power measures, as evaluated in the cohort of high-energy physicists. For comparison purposes, an alternative measure of diversity based on Stirling’s index [56], with prior applications to studies of interdisciplinarity [57, 58] is evaluated [^17] The Stirling-based diversity measured is evaluated as: $D_{\text{Stirling}}=1-\sum_{k,k^{\prime}}d_{kk^{\prime}}I_{ak}I_{ak^{\prime}}$ (15) Where $d_{kk^{\prime}}$ is the fraction of scientists who have more expertise than average in both $k$ and $k^{\prime}$ among those that have expertise in one or the other (i.e., a similarity matrix); following previous approaches for measuring research interdisciplinarity [57, 58].. A measure of brokerage is also considered [^18] We evaluated brokerage as the amount of pairs of scientists that have collaborated with a given physicist while having no common collaborator except for this physicist. This effectively measures the extent to which this physicist connects otherwise disconnected scientists..
@@ A.6 Learning costs and optimal transport
The MCMC algorithm from [52] is run on 500,000 iterations of the “MetroMC” algorithm, using a “P1” prior (that is, a prior such that $\sum_{kk^{\prime}}C_{kk^{\prime}}=C_{0}=\mathrm{cst}$). More precisely, we assume that:
$P(c_{kk^{\prime}}|p_{kk^{\prime}})$ $=\dfrac{1}{Z}\dfrac{1}{\prod_{kk^{\prime}}c_{kk^{\prime}}^{1/2}}\exp{\left(-\alpha D_{KL}(c_{kk^{\prime}}||p_{kk^{\prime}})\right)}\text{ with}c_{kk^{\prime}}=C_{kk^{\prime}}/C_{0}$ (16)
[…]
The second equation is sometimes referred to as the entropic prior [59, 60]. The mean posterior values of $C_{k,k^{\prime}}$ are shown in Figure 14(a), as a function of the knowledge gap from $k$ to $k^{\prime}$. The knowledge gap $1-\nu_{kk^{\prime}}$ is the fraction of experts in $k$ who do not hold significant expertise in $k^{\prime}$ ($\nu$ is shown in Figure 14(b)). A significant correlation is found ($R=-0.72$). This is true also for the replication dataset of computational linguistics research (Appendix A.9), for which $R=-0.63$.
Figure 14: Cost matrix $C_{kk^{\prime}}$. The cost matrix represents the cost of shifting a unit of cohort’s research attention from $k$ to $k^{\prime}$.
(a) Cost of shifting a unit of cohort’s research attention from $k$ to $k^{\prime}$ as a function of the fraction of physicists with expertise in $k^{\prime}$ among those with expertise in $k$ ($\nu_{kk^{\prime}}$).
(b) Fraction $\nu_{kk^{\prime}}$ of physicists with expertise in the horizontal research area ($k$) who also hold expertise in the vertical research area ($k^{\prime}$).
Figure 15: Learning costs and knowledge gap. The cost of shifting attention from one area to another (left plot) is in great part determined by how frequently individuals hold knowledge in both areas (right plot).
@@ A.9 Replication corpus
For purposes of replication, certain analyses have been reproduced on the ACL anthology corpus of Computational Linguistics research [15]. A comparison with the HEP corpus in shown in Table 7. The ACL corpus is smaller, presumably due to a lack of exhaustivity. The requirement on the quantity of publications has been lowered (three per time period instead of five) in order to maintain a reasonable size of cohort. The choice of time periods is arbitrary (i.e. it was not informed by the context). Only 14 topics were deemed to carry scientifically meaningful content upon manual inspection.